# Speech Foundation Models Generalize to Time Series Tasks from Wearable Sensor Data

### v2

@@ 1 Introduction and Related Work
Time series models have been trained for applications spanning numerous domains—including health, activity recognition, gesture recognition, weather forecasting, and infrastructure modeling. Classical time series modeling methods include dynamic time warping, shapelet-based methods, convolution-based methods like ROCKET, and numerous other approaches (Bagnall et al. (2017); Middlehurst et al. (2024)). Recent work has also explored deep learning-based methods and representation learning strategies to enable shared learning among tasks (Xu et al. (2023, 2024); Abbaspourazad et al. (2023); Erturk et al. (2025)). In the realm of physiological and wearable sensor data, most approaches have focused on within-domain representation learning where there is a domain match between pre-training data and evaluation tasks. Prior works on transfer learning for time series modeling often focus on time series data like power consumption, traffic, and weather and often on forecasting tasks (Woo et al. (2024); Liu et al. (2023); Jin et al. (2023)).
Voice2Series (Yang et al. (2021)) explored re-programming speech processing models for time series tasks using task-specific target data along with a transformer-based speech model, and found strong performance across the evaluated UCR datasets—including some sensor-based tasks like ECG modeling. Voice2Series trained speech embedding models from scratch, re-trained layers within the model for each time series task, and then identified source-to-target label maps for inference in the new domains. While successful, this approach required both re-training and label mapping, as well as from-scratch training of speech foundation models. Between domain knowledge transfer has been successful in other areas—for instance, the Audio Spectrogram Transformer (Gong et al. (2021)) showed improved performance on speech tasks by training a ViT model on ImageNet data.
Figure 1: Time-series data from modalities such as audio, accelerometer (Accel.), electrocardiogram (ECG), and photoplethysmogram (PPG) signals contain rich temporal and spectral characteristics, including frequency band powers, periodic patterns, and distinctive waveform shapelets.
@@ 2 Methods and Results
Activity Classification We evaluated activity classification on four 3-axis accelerometer data datasets spanning three sensor positions and two window sizes. In both evaluations, we included a benchmark using a Random Forest and engineered features, as in Yuan et al. (2024). We included results and benchmarks from two window sizes along with performance comparisons from Xu et al. (2024); Yuan et al. (2024); Haresamudram et al. (2022): (1) following Yuan et al. (2024), a leave one subject out subject out evaluation scheme with 10 second windows at 100 Hz with the PAMAP2 wrist data Reiss and Stricker (2012) (8 classes, n=2,869) and the Opportunity dataset (4 classes, n=3,842) Roggen et al. (2010), and (2) following Haresamudram et al. (2022), a 5-fold cross-validation evaluation on 2 second windows of 100 Hz 3-axis accelerometer data: HHAR Reyes-Ortiz et al. (2015) wrist data (6 classes, n=3,370), Motionsense waist data (6 classes, n=14,121), Malekzadeh et al. (2018), and PAMAP2 leg data (12 classes, n=9,709). The 2 second windows were upsampled by a factor of 2 to have sufficient sample lengths to use with the pre-trained speech models. The embeddings extracted from 10 second windows were pooled across the time dimension before being passed to the probe, in order to reduce dimensionality. Table 1 shows results from five activity classification evaluations: two with 10 second windows following the evaluation in Yuan et al. (2024) and three with 2 second windows following the evaluation in Haresamudram et al. (2022).
Arrhythmia Detection (ECG) We conducted binary arrhythmia classification using the MIT-BIH dataset Moody (1983), using 10 second windows sampled at 250 Hz. The extracted embeddings were pooled across the time dimension before being passed to the probe, in order to reduce dimensionality. We compare to previous results on this dataset reported in Xu et al. (2023), including a baseline supervised model and self-supervised models trained on ECG data. Results are reported in Table 2.
@@
Yuan et al. (2024) Eval Haresamudram et al. (2022) Eval
@@
Table 1: Activity classification results. Probes with pre-trained speech models compared to probes with pre-trained IMU models from Yuan et al. (2024) "(Yuan’24)" and Haresamudram et al. (2022) "(Hare’22)" and a baseline Random forest model with engineered features.
Mood Classification (PPG) We conducted four-class mood classification (baseline, stress, amusement, and meditation) using PPG data from WESAD Schmidt et al. (2018), in one minute windows sample at 64 Hz. Results are reported in Table 2, along with comparative results with identical evaluation from Xu et al. (2023) including a baseline supervised model and self-supervised models trained on PPG data.
Arrythmia Detection (ECG) Mood Classification (PPG)
@@
Table 2: Arrhythmia detection results and mood classification results. Probes with pre-trained speech models compared to probes with pre-trained ECG models and pre-trained PPG models from Xu et al. (2023).
@@ 3 Discussion
Probes trained on pre-trained speech foundation models had the best or competitive performance across tasks when using MLP probes trained on embeddings extracted from early layers. Pre-trained speech representations outperformed both baselines and self-supervised models trained directly on sensor data for most tasks—activity classification with two second windows, mood classification, and arrhythmia detection. Probes trained with pre-trained speech representations had competitive performance for activity classification with 10 second windows, though lower scores than the RelCon model trained on accelerometer data.
@@
4 Conclusions and Future Work
The analyses suggest that learned speech feature extractors can be particularly effective for wearable sensor tasks. Models like HuBERT and wav2vec 2.0 benefit from large, well-curated speech datasets that enable the learning of rich and transferable representations. In contrast, wearable sensor datasets are typically smaller and task-specific. Our results show that cross-domain learning—leveraging speech data—improves performance on data-scarce tasks involving wearable sensors.
While speech and the investigated sensor domains (accelerometer data, ECG, and PPG) share enough commonalities to enable zero-shot transfer, the investigated sensor data streams were sampled at lower frequencies (100 Hz, 250 Hz, and 64 Hz respectively) than speech data. The pre-trained foundation models were designed for speech input at 16,000 Hz. Training strategies for foundation models that better enable learning over both the low and high frequency space could help improve performance across modalities and will also be explored in future work.

### v3

@@ 1 Introduction and Related Work
Time series models have been trained for applications spanning numerous domains—including health, activity recognition, gesture recognition, weather forecasting, and infrastructure modeling. Classical time series modeling methods include dynamic time warping, shapelet-based methods, convolution-based methods like ROCKET, and numerous other approaches (bagnall2017great; middlehurst2024bake). Recent work has also explored deep learning-based methods and representation learning strategies to enable shared learning among tasks (xu2023rebar; xu2024relcon; abbaspourazad2023large; erturk2025beyond). In the realm of physiological and wearable sensor data, most approaches have focused on within-domain representation learning where there is a domain match between pre-training data and evaluation tasks. Prior works on transfer learning for time series modeling often focus on time series data like power consumption, traffic, and weather and often on forecasting tasks (woo2024unified; liu2023itransformer; jin2023time).
Voice2Series (yang2021voice2series) explored re-programming speech processing models for time series tasks using task-specific target data along with a transformer-based speech model, and found strong performance across the evaluated UCR datasets—including some sensor-based tasks like ECG modeling. Voice2Series trained custom speech embedding models from scratch, re-trained layers within the model for each time series task, and then identified source-to-target label maps for inference in the new domains. While successful, this approach required both re-training and label mapping, as well as from-scratch training of speech foundation models. We use publicly available frozen pre-trained models directly as feature extractors, and train only lightweight probes for each task, without any re-training the embedding model backbone or label mapping. Similar between domain knowledge transfer has been successful in other areas—for instance, the Audio Spectrogram Transformer (gong2021ast) showed improved performance on speech tasks by training a ViT model on ImageNet data.
Figure 1: Time-series data from modalities such as speech, accelerometer (Accel.), electrocardiogram (ECG), and photoplethysmogram (PPG) signals contain rich temporal and spectral characteristics, including frequency band powers, periodic patterns, and distinctive waveform shapelets.
@@ 2 Methods and Results
Activity Classification We evaluated activity classification on four 3-axis accelerometer data datasets spanning three sensor positions and two window sizes. In both evaluations, we included a benchmark using a Random Forest and engineered features, as in yuan2024self. We included results and benchmarks from two window sizes along with performance comparisons from xu2024relcon; yuan2024self; haresamudram2022assessing: (1) following yuan2024self, a leave one subject out subject out evaluation scheme with 10 second windows at 100 Hz with the PAMAP2 wrist data reiss2012introducing (8 classes, n=2,869) and the Opportunity dataset (4 classes, n=3,842) roggen2010collecting, and (2) following haresamudram2022assessing, a 5-fold cross-validation evaluation on 2 second windows of 100 Hz 3-axis accelerometer data: HHAR reyes2015smartphone wrist data (6 classes, n=3,370), Motionsense waist data (6 classes, n=14,121), Malekzadeh:2018:PSD:3195258.3195260, and PAMAP2 leg data (12 classes, n=9,709). The 2 second windows were upsampled by a factor of 2 to have sufficient sample lengths to use with the pre-trained speech models. The embeddings extracted from 10 second windows were pooled across the time dimension before being passed to the probe, in order to reduce dimensionality. Table 1 shows results from five activity classification evaluations: two with 10 second windows following the evaluation in yuan2024self and three with 2 second windows following the evaluation in haresamudram2022assessing.
Arrhythmia Detection (ECG) We conducted binary arrhythmia classification using the MIT-BIH dataset moody1983new, using 10 second windows sampled at 250 Hz. The extracted embeddings were pooled across the time dimension before being passed to the probe, in order to reduce dimensionality. We compare to previous results on this dataset reported in xu2023rebar, including a baseline supervised model and self-supervised models trained on ECG data. Results are reported in Table 2.
@@
yuan2024self Eval haresamudram2022assessing Eval
@@
Table 1: Activity classification results. Probes with pre-trained speech models compared to probes with pre-trained IMU models from yuan2024self "(Yuan’24)" and haresamudram2022assessing "(Hare’22)" and a baseline Random forest model with engineered features.
Mood Classification (PPG) We conducted four-class mood classification (baseline, stress, amusement, and meditation) using PPG data from WESAD schmidt2018introducing, in one minute windows sample at 64 Hz. Results are reported in Table 2, along with comparative results with identical evaluation from xu2023rebar including a baseline supervised model and self-supervised models trained on PPG data.
Arrhythmia Detection (ECG) Mood Classification (PPG)
@@
Table 2: Arrhythmia detection results and mood classification results. Probes with pre-trained speech models compared to probes with pre-trained ECG models and pre-trained PPG models from xu2023rebar.
@@ 3 Discussion
Model performance Probes trained on pre-trained speech foundation models had the best or competitive performance across tasks when using MLP probes trained on embeddings extracted from early layers. Pre-trained speech representations outperformed both baselines and self-supervised models trained directly on sensor data for most tasks—activity classification with two second windows, mood classification, and arrhythmia detection. Probes trained with pre-trained speech representations had competitive performance for activity classification with 10 second windows, though lower scores than the RelCon model trained on accelerometer data. The longer windows may have had enough information to better leverage the learned motifs in RelCon – better understanding how modeling strategy and sample window intersect will be explored in the future.
@@
Limitations While speech and the investigated sensor domains (accelerometer data, ECG, and PPG) share enough commonalities to enable shared enocoders, the investigated sensor data streams were sampled at lower frequencies (100 Hz, 250 Hz, and 64 Hz respectively) than speech data (typically at 16,000 Hz with these models). Future experiments will include ablations to assess impact of sampling rates, input processing, and window lengths as well as training strategies to better enable learning across diverse sample frequencies. We present, to our knowledge, a set of first experiments showing that speech models generalize to wearable time series tasks – additional validation is needed to further develop and understand how this methodology might be deployed, like evaluation with additional datasets, probing modeling biases, and further investigating interpretability.
Conclusions The analyses show that pre-trained speech models can be effective feature extractors for wearable sensor tasks. Models like HuBERT and wav2vec 2.0 are trained with large, well-curated speech datasets, learning of rich and transferable representations. In contrast, wearable sensor datasets are typically smaller and task-specific. The convolution encoders from speech models were particularly impactful, suggesting that motif-learning in time series models may be particularly impactful. Our results show that cross-domain learning—leveraging speech data—has strong performance on data-scarce tasks involving wearable sensors, suggesting cross-modality adaption and learning as an impactful modeling strategy.
4 Appendix
4.1 Datasets
PAMAP2 (Wrist) We extracted 10 second segments at 100 Hz from PAMAP2 reiss2012introducing, and used a leave one subject out evaluation scheme following the procedure in yuan2024self. The evaluation included eight classes: lying, sitting, standing, walking, ascending stairs, descending stairs, vacuum cleaning, and ironing. The experiments included n=2,860 samples from eight participants.
Opportunity (Wrist) We extracted 10 second segments at 100 Hz from roggen2010collecting, and used a leave one subject out evaluation scheme following the procedure in yuan2024self. The evaluation included four classes: sitting, standing, walking, lying down. The experiments included n=3,842 samples from four participants.
HHAR (Wrist) We extracted 2 second segments at 100 Hz from reyes2015smartphone, and used a 5-fold cross validation evaluation scheme using the procedure in haresamudram2022assessing. The evaluation included six classes: stairs down, stairs up, walk, bike, stand, sit. The experiments included n=3,370 samples.
Motionsense (Waist) We extracted 2 second segments at 100 Hz from Malekzadeh:2018:PSD:3195258.3195260, and used a 5-fold cross validation evaluation scheme using the procedure in haresamudram2022assessing. The evaluation included six classes: stairs down, stairs up, walk, jog, stand, sit. The experiments included n=14,121 samples.
PAMAP2 (Leg) We extracted 2 second segments at 100 Hz from reiss2012introducing, and used a 5-fold cross validation evaluation scheme using the procedure in haresamudram2022assessing. The evaluation included twelve classes: rope jumping, lying, sitting, standing, walking, running, cycling, Nordic walking, ascending stairs, descending stairs, vacuum cleaning, ironing. The experiments included n=9,709 samples.
MIT-BIH We extracted 10 second segments at 250 Hz from moody1983new. We followed the pre-processing and evaluation process as described in xu2023rebar to allow for comparison with prior results from self-supervised models trained in-domain.
WESAD We extracted 60 second segments at 64 Hz from schmidt2018introducing. We followed the pre-processing and evaluation process as described in xu2023rebar to allow for comparison with prior results from self-supervised models trained in-domain.